OpenStudy (anonymous):

Related rates problem: The beacon on a lighthouse makes one revolution every 20 seconds. The beacon is 300 feet from the nearest point, P, on a straight shoreline. Find the rate at which the ray of light moves along the shore at a point 200 feet from P. (In other words, how fast would you have to drive a car along the shoreline so that you keep up with the light constantly if you were 200 feet from P on the shoreline).

OpenStudy (anonymous):



OpenStudy (anonymous):

tan(T) = 200/300 dT/dt = pi/10 radians per second

OpenStudy (anonymous):

10 = (300(dx/dt) - 200(dy/dt))/(300^2) , do these look correct so far?

OpenStudy (anonymous):

http://www.jiskha.com/display.cgi?id=1339826996 for those interested